# CoSteer: Collaborative Decoding-Time Personalization via Local Delta Steering

### v1

@@ B.4 Prompts
We use exactly the same prompt as Cogenesis [35] to have the gpt-4o evaluate the generated content. The specific prompts are shown in 6 and 7.
Appendix C Limitations and broader impact
C.1 Limitations and future work
While CoSteer balances privacy and quality effectively, its performance depends on the SLM’s ability to capture meaningful personalization signals, which may falter for highly complex user contexts. Future directions include exploring lightweight SLMs optimized for personalization, extending the framework to multi-modal tasks, and investigating federated learning paradigms to further enhance privacy and scalability. By addressing these challenges, CoSteer paves the way for adaptive, user-centric AI systems that harmonize global knowledge with localized individuality.
C.2 Broader impact
The CoSteer framework presents transformative potential for deploying personalized AI systems in privacy-sensitive and resource-constrained environments. Below, we discuss its broader societal implications, including both positive contributions and potential risks.
Positive impacts:
  • Enhanced privacy preservation: By processing sensitive user data locally and transmitting only privacy-preserving steering signals, CoSteer mitigates risks of data breaches and misuse. This aligns with growing global demands for ethical AI that respects user confidentiality.
  • Democratization of personalization: CoSteer enables high-quality personalized experiences even on low-resource devices, bridging the gap between cloud-scale AI capabilities and edge computing. This could empower users in regions with limited connectivity or computational infrastructure.
  • Efficient resource utilization: The framework reduces reliance on constant cloud computation, lowering energy consumption and operational costs for large-scale deployments, contributing to sustainable AI practices.
  • Applications in critical domains: CoSteer’s privacy-aware design makes it suitable for healthcare, education, and financial services, where personalized insights must balance utility with strict data protection.
Potential risks and mitigations:
  • Misuse of personalization: Malicious actors could exploit the framework to generate targeted misinformation or manipulative content. Safeguards like user-controlled steering parameters and content moderation mechanisms are essential.
  • Bias amplification: Localized personalization may inadvertently reinforce user-specific biases if SLMs learn from skewed historical data[51]. Regular audits and fairness-aware training for SLMs could address this.
  • Dependence on local models: Performance hinges on the SLM’s ability to capture nuanced personal contexts. Future work should focus on lightweight, robust SLMs to prevent quality degradation in complex scenarios.
  • Accessibility disparities: While edge-friendly, CoSteer assumes access to modern devices. Partnerships with policymakers and NGOs could ensure equitable access to such technologies across socioeconomic groups.
@@ C.2 Broader impact
Cogenesis [Task] Design an invitation for ’Homes for Humanity’ charity event: Utilize your bi-monthly volunteering work and personal charm to craft an invitation. [User Profile] Age: 57 years Name: Martin Reynolds Occupation: Senior Real Estate Agent Location: Charlotte, North Carolina Personal traits: Detail-oriented; Personable; Strategic thinker; Avid golfer; Enjoys weekend DIY projects Writing style: Clear and concise; Persuasive and sales-driven; Friendly, often includes anecdotes related to golf or DIY projects; Professional, with a touch of personal charm Privacy info: Successfully renegotiated the lease terms for the company office last month; Celebrated 30th wedding anniversary with a surprise garden party; Volunteers bi-monthly at the local ’Homes for Humanity’ charity; Recently started attending a beginner’s pottery class on Sundays; Has a standing golf game every Wednesday afternoon with industry peers; Is secretly learning Spanish online to communicate with more clients; Adopts a new 2̆018Real Estate Investment2̆019 tip of the month for social media Smart device usage: Email: Sent a client proposal at 8:45 AM - ’184 Maple Drive Listing Proposal’; Calendar: Golf game scheduled for May 5, 1:30 PM with note - ’Meet Greg at Fairways Club’; Photos: Pictures of a kitchen renovation ’Before and After’ at 3:52 PM; App: Logged into ’HouseFlipper’ real estate investment app at 9:17 PM; Text Message: Received at 2:15 PM - ’Happy Anniversary! Table booked at La Trattoria.’; Voice Assistant: Asked for ’directions to the nearest hardware store’ at 10:10 AM; Purchase: Ordered a ’Stainless Steel Faucet’ from ’DIYHome’ app at 7:45 PM; Notification: ’Your Spanish lesson starts in 5 minutes’ at 7:55 AM; Search: Googled ’Energy-efficient home improvements’ at 10:25 AM Ai assistant usage: Composing a visually rich monthly newsletter for property investors; Creating client-specific follow-up 2̆00bemails capturing unique details of each viewing; Generating personalized invitations for exclusive open houses and real estate events" [writing history] [query]:Draft a proposal for ’184 Maple Drive’ listing: Using the information from your morning email, flesh out a comprehensive proposal for the ’184 Maple Drive’ property. [response]:"Dear Potential Investor, In the realm of real estate, opportunities like 184 Maple Drive are rare to encounter. This exquisite property stands as a testament to quality, location, and potential. Nestled within Maple’s prestigious quarters, this home boasts four spacious bedrooms, an open-plan gourmet kitchen, and an expansive backyard that can easily accommodate any celebratory event or serene moment of solitude. I remember how just last year, two blocks over, I closed a deal on a similar property2̆014a testament to the rapidly appreciating value in this neighborhood. Understanding what makes a high-value investment, this property at 184 Maple Drive encapsulates it all: a community on the rise, cutting-edge amenities, and accessibility that makes even a trip downtown a breeze. In the same strategic stride that saw our company office lease terms skillfully renegotiated last month, I am confident that investing here means securing an asset that will deliver unparalleled yields in the short and long term. As a strategic thinker, I’ve always marveled at homes that offer more than just shelter, and this property is the embodiment of strategic living2̆014it2̆019s where energy meets efficiency, embracing the trend for greener living. Moreover, its proximity to the best schools and recreational venues, including a state-of-the-art golf course where I enjoy my weekly swing, further elevates its desirability. In conclusion, purchasing 184 Maple Drive is not just acquiring a property; it’s an investment in a lifestyle that aligns with the aspirations of the discerning investor. Warm Regards, Martin Reynolds Senior Real Estate Agent, Queen City Realty"
@@
Overall Assesment [Instruction] Please act as an impartial evaluator and assess the quality of the AI assistant’s response to the user question shown below. Your assessment should focus on how well the response aligns with the user’s personalized profile and writing history. Evaluate factors such as the response’s adherence to the user’s personal style, consistency with their profile, helpfulness, relevance, accuracy, depth, creativity, and level of detail. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: [[rating]], for example: Rating: [[5]]. [User Profile and Writing History] <profile_info> <writing_history> [Question] <question> [The Start of Assistant’s Answer] <answer> [The End of Assistant’s Answer]
@@
Personal Assessment [Instruction] Please act as an impartial judge an evaluate the AI assistant’s response based on its alignment with the user’s personal profile and writing history. Focus your assessment on the personalization aspects of the response, including its adherence to the user’s unique style, preferences, and consistency with their profile. Consider how well the response addresses the user’s individual needs and interests. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: [[rating]], for example: Rating: [[5]]. [User Profile and Writing History] <profile_info> <writing_history> [Question] <question> [The Start of Assistant’s Answer] <answer> [The End of Assistant’s Answer]

### v2

@@ B.7 Prompts for LLM-as-a-Judge
We use exactly the same prompt as Cogenesis (Zhang et al., 2024b) to have the gpt-4o evaluate the generated content. The specific prompts are shown in Figure 6 and Figure 7.
B.8 Baseline Implementation Details
We compare our proposed method against three representative baseline approaches:
CoS / LA
Context Steering (CoS) (He et al., 2025) and Linear Alignment (LA) (Gao et al., 2024) are conceptually equivalent methods. They steer the generation of a base model by modifying its output log-probabilities. The final log-probability for a token $x_{i}$ is calculated by adding a scaled difference term to the base model’s prediction, which is conditioned on no context ($\emptyset$). The formulation is as follows:
$\log P^{\prime}(x_{i})=\log P(x_{i}|\emptyset,\mathcal{P})+\lambda\cdot(\log P(x_{i}|C,\mathcal{P})-\log P(x_{i}|\emptyset,\mathcal{P}))$
where $\mathcal{P}$ represents the base model parameters, $C$ is the provided context, and $\lambda$ is a scaling hyperparameter. Following the implementations in the original works, we searched for the optimal $\lambda$ from the set $\{1.5,2.0,2.5\}$ and report the best-performing result for each task.
Supervised Fine-Tuning (SFT)
To create a strong task-specific baseline, we fine-tuned the small language model using Low-Rank Adaptation (LoRA). For each of our three datasets, we randomly sampled 1,000 data points for training. The model was then trained for three epochs. Key training parameters included a learning rate of $1\times 10^{-4}$ with a cosine scheduler, a maximum sequence length of 1024, a weight decay of 0.05, and bf16 precision for efficiency. We utilized packing to handle variable-length sequences effectively.
Proxy-Tuning
This method (Liu et al., 2024) enhances a large base model by incorporating signals from a smaller, specialized proxy model. The final probability distribution is calculated by adjusting the logits of the large model. We adapt the original formula to better reflect the roles of the different models in our setup:
$p(x_{t}|x_{<t})=\text{softmax}[s_{\text{LLM}}(x_{t}|x_{<t})+s_{\text{SLM-SFT}}(x_{t}|x_{<t})-s_{\text{SLM-base}}(x_{t}|x_{<t})]$
Here, $s_{\text{LLM}}$ represents the logits from the large, pre-trained language model. The term $s_{\text{SLM-SFT}}$ refers to the logits from the small language model that has been fine-tuned (the SFT baseline), and $s_{\text{SLM-base}}$ refers to the logits from the original, pre-trained small language model.
B.9 Parameter Sensitivity Analysis
Results are shown in Figure 2. Below, we analyze each parameter.
B.10 Robostness to Noise
To supplement the discussion in Section 5.1, this section provides the detailed experimental setup and result analysis for our robustness tests.
B.10.1 Robustness to Realistic Noise (Weaker Retriever)
Experimental Setup
In many real-world systems, the retrieval component may not be state-of-the-art. To simulate this “realistic noise”, where retrieved context is topically relevant but potentially less precise, we replaced the powerful bge-m3 retriever with the classic, non-neural BM25 retriever. This setup tests whether CoSteer is overly dependent on a high-quality retriever or if it can adapt to a more common, weaker signal. We use Qwen7B-1.5B pair to conduct the experiment.
Results and Analysis
The results are presented in Table 9. Even with the noisier context provided by BM25, CoSteer demonstrates strong performance, achieving the second-best results across nearly all metrics, only behind the full LLM w/ oracle. For instance, in the Abstract generation task, CoSteer (43.08 R-1) significantly outperforms both the SLM w/ (39.51 R-1) and the strong LLM w/o (39.81 R-1) baselines. This confirms that our framework is not brittle and can effectively leverage imperfect but relevant local information, a critical capability for practical deployment.
Table 9: Robustness to a weaker retriever (BM25). Even with this noisier context, CoSteer consistently outperforms both the SLM w/ and the LLM w/o baselines, demonstrating its robustness in practical scenarios.
Abstract Review Writing
R-1 R-L MET R-1 R-L MET R-1 R-L MET
SLM w/o 36.48 17.74 27.57 20.40 10.39 10.39 25.11 11.09 17.51
SLM w/ 39.51 21.92 27.74 23.42 12.00 13.14 24.15 12.32 18.22
LLM w/o 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49
CoSteer 43.08 23.49 28.08 31.95 15.36 19.45 25.47 12.46 22.83
LLM w/ 44.17 24.55 30.83 33.70 15.42 22.47 28.90 13.89 23.84
B.10.2 Robustness to Purely Irrelevant (Adversarial) Noise
Experimental Setup
To push the limits of our framework, we designed a more adversarial scenario to test a potential failure mode: providing the model with context that is completely irrelevant to the task. For example, when generating an Abstract, we supplied context examples drawn from the Review task. This setup evaluates whether the model can ignore distracting, irrelevant information or if it gets catastrophically misled. We use Qwen7B-1.5B pair to conduct the experiment.
Results and Analysis
Table 10 shows the outcomes of this stress test. As expected, the performance of all models degrades compared to using relevant context. However, the key observation is in the *manner* of degradation. CoSteer (e.g., 39.59 R-1 on Abstract) experiences a controlled performance drop but still comfortably outperforms the specialized SLM w/ (34.91 R-1). It does not collapse. This graceful degradation suggests that the strong inductive bias from the frozen base LLM acts as a safeguard, preventing the steering mechanism from being entirely derailed by nonsensical context. The model learns to rely more on its pretrained knowledge when the provided context is useless, showcasing a desirable level of robustness.
Table 10: Robustness to purely irrelevant (adversarial) noise. We tested a failure mode by providing irrelevant context (e.g., Review examples for the Abstract task). The framework degrades gracefully rather than failing catastrophically, still outperforming the local SLM.
Abstract Review Writing
R-1 R-L MET R-1 R-L MET R-1 R-L MET
SLM w/o 36.48 17.74 27.57 20.40 10.39 10.39 25.11 11.09 17.51
SLM w/ 34.91 18.31 24.65 22.15 11.56 12.25 22.68 9.71 18.03
LLM w/o 39.81 20.53 25.56 30.15 14.04 17.71 27.64 11.93 21.49
CoSteer 39.59 21.17 24.16 30.45 14.11 17.76 24.71 10.80 20.19
LLM w/ 38.43 20.72 23.65 31.69 14.49 21.01 25.48 11.22 20.65
B.11 Results and Implementation of Cross-Architecture Collaboration
Table 11: Performance of CoSteer in a cross-architecture setting, using Llama-3.1-8B as the LLM and Qwen2.5-1.5B as the SLM. We compare two vocabulary-agnostic strategies: vocabulary mapping (CoSteer_map) and a more universal byte-level fusion (CoSteer_byte). The results show that both methods enable effective collaboration, validating our framework’s generalization capability across different model families.
Abstract Review Writing
R-1 R-L MET R-1 R-L MET R-1 R-L MET
SLM w/o 36.48 17.64 27.57 20.40 10.39 10.39 25.21 11.09 17.51
SLM w/ 39.75 22.03 27.76 23.08 12.40 12.94 22.89 11.46 17.12
LLM w/o 39.33 20.69 29.41 34.58 15.32 22.10 26.93 12.35 21.82
CoSteer_map 41.82 23.40 31.79 33.72 15.63 23.22 25.01 11.78 21.10
CoSteer_byte 42.84 22.69 32.03 33.58 15.24 23.50 23.29 10.94 20.14
LLM w/ 43.91 23.93 32.01 36.39 15.95 23.56 30.54 14.02 23.81
The experiments were conducted using Llama-3.1-8B as the LLM and Qwen2.5-1.5B as the SLM. Results are shown in Table 11. Below we details the two strategies we implemented to facilitate collaboration between models with different architectures and tokenizers, as discussed in Section 5.2.
B.11.1 Vocabulary Mapping (CoSteer_map)
Principle
The vocabulary mapping approach establishes a shared communication channel by restricting fusion to the intersection of the two models’ vocabularies, allowing agreement on a common set of semantic units.
Implementation
Our implementation follows a three-step process at each generation step:
  1. 1.
    Vocabulary Intersection: Before generation begins, we create a mapping between the two models. We extract the token strings from both tokenizers and find their intersection. Two tensors, llm_intersect_ids and slm_intersect_ids, are then created to store the corresponding token IDs for this shared vocabulary, ensuring a deterministic alignment.
  2. 2.
    Logit Projection: At each decoding step, we obtain the native logit outputs from both the LLM and SLM. Using torch.index_select, we project these full-vocabulary logits onto the smaller, shared vocabulary space defined by the intersection.
  3. 3.
    Optimization and Remapping: The CoSteer optimization is applied in this shared space. The resulting token is mapped back to each model’s native ID space via the pre-computed tensors and appended to their respective input sequences.
While conceptually simple and effective, a key limitation of this approach is that the intersection can be small for models with dissimilar tokenizers, potentially limiting the expressivity of the generation.
B.11.2 Byte-level Fusion (CoSteer_byte)
Principle
To overcome the limitations of vocabulary mapping, CoSteer_byte operates entirely in a shared byte space. Instead of aligning token IDs, it projects each model’s token-level logit distribution into a common 257-dimensional space (Hayase et al., 2025): 256 dimensions for possible next bytes (0–255) and one special “commit” dimension (byte 256) indicating that the current token is complete. This enables direct byte-level probabilistic fusion of LLM and SLM outputs.
Implementation
The byte-level strategy integrates with our optimization as follows:
  1. 1.
    Byte Projection: For each model, we map its token-level logits to a byte-level log-probability distribution over the 257-dimensional space. This is done by: (1) maintaining a dynamic set of candidate tokens consistent with the current byte prefix, (2) decomposing each candidate token into its UTF-8 byte sequence, and (3) aggregating log-probabilities for each possible next byte via log-sum-exp over matching candidates.
  2. 2.
    Optimization and Byte Sampling: CoSteer fuses the three byte-level distributions into a single policy, from which the next byte is sampled. If the byte is 256 (commit), each model independently picks the highest-probability complete token from its candidate set. Despite differing token IDs due to tokenizer mismatch, these tokens decode to the same UTF-8 string, ensuring semantic alignment. Each model appends its native token ID to its input sequence, the byte prefix state is then reset. Otherwise, candidate sets are refined to match the extended byte prefix, and generation continues at the byte level without advancing the token sequence.
This approach requires no vocabulary overlap and supports arbitrary model pairs. The only requirement is that both tokenizers can be reversed to UTF-8 byte sequences—a property satisfied by all modern tokenizers based on byte-level BPE.
B.12 Implementation and Analysis of CoSteer Variants
This section provides a detailed description of the LightCoSteer and AdaCoSteer variants, followed by a comprehensive analysis of their performance and efficiency trade-offs.
B.12.1 LightCoSteer: Single-Step Steering
Motivation and Implementation
The primary source of computational overhead in the full CoSteer framework is the iterative optimization process within the FTRL algorithm. LightCoSteer is designed to eliminate this overhead entirely. We achieve this by setting the number of iterations to one ($T=1$). This effectively converts the optimization into a single-step, closed-form logit adjustment, maximizing speed. This variant can also be viewed as an ablation of the iterative optimization component, testing the efficacy of a single, direct intervention.
Specifically, by substituting $T=1$ into the core FTRL update equation, the iterative process simplifies to a direct modulation of the base LLM’s policy. The resulting policy for selecting an action (token) $a$ is equivalent to:
$\pi(a)\propto\exp\left(\log\pi_{0}(a)+\beta^{*}\left(\log\pi^{*}_{\text{pers}}(a)-\log\pi^{*}_{\text{base}}(a)\right)\right)$ (17)
where $\pi_{0}$ is the base LLM policy, $\pi^{*}_{\text{pers}}$ and $\pi^{*}_{\text{base}}$ are the SLM policies with and without context respectively, and $\beta^{*}$ is a weighting hyperparameter.
B.12.2 AdaCoSteer: Adaptive Steering
Motivation and Implementation
While LightCoSteer addresses computational load, it does not reduce the number of communication rounds between the local SLM and the remote LLM, as every token still requires steering. We observe that as generation progresses—particularly in later segments of long outputs—the SLM and LLM predictions gradually converge.
Motivated by this, AdaCoSteer implements an adaptive strategy to reduce unnecessary communication. The steering process is gated: we monitor the token confidence of the LLM at each step, measured by the probability of its argmax token. If this confidence score exceeds a predefined threshold $\tau$ for $k$ consecutive steps, we deactivate CoSteer. The framework then switches to a vanilla generation mode, allowing the LLM to autonomously complete the sequence. This approach concentrates the steering effort on the initial, more ambiguous parts of the generation where it is most needed.
B.12.3 Performance
We present the performance and efficiency results of the two variants compared to the full CoSteer framework in Table 12, Table 14, and Table 14.
Table 12: Performance comparison of different CoSteer framework variants.
Abstract Review Writing
R-1 R-L MET R-1 R-L MET R-1 R-L MET
AdaCosteer 40.97 22.04 25.96 29.94 14.22 17.15 26.48 12.12 22.24
LightCosteer 41.79 21.86 26.44 31.46 14.80 18.31 27.33 12.19 21.89
CoSteer 42.98 23.61 28.20 32.72 15.92 20.36 25.93 12.38 22.84
Table 13: Inference speed and time complexity
analysis for CoSteer and its variants.
Method Time Complexity Speed (tok/s)
Vanilla Gen. $L(n)$ 23.88
AdaCosteer $L(n)+C(t_{c})+T(c)$ 20.65
LightCosteer $L(n)+C(n)+T(n)$ 13.73
CoSteer $L(n)+C(t_{n})+T(n)$ 9.44
Table 14: Computational cost (TFLOPs) as a function of context length (C).
Method C=10 C=100 C=1000
LLM w/o 0.87 0.87 0.87
LLM w/ 0.96 1.74 9.98
CoSteer 1.21 1.35 2.98
Effectiveness Analysis
Table 12 shows the performance trade-offs. The full CoSteer framework consistently achieves the highest scores across all tasks. LightCoSteer follows closely, demonstrating that a single-step adjustment retains a significant portion of the framework’s benefits. AdaCoSteer experiences a slightly larger performance drop, which is expected as it deliberately stops steering in later, potentially less critical, generation stages.
B.12.4 Efficiency and Overhead Analysis
Empirical Results
Table 14 quantifies the throughput (tokens/second) of our proposed variants, while Table 14 compares their computational cost (FLOPS) against baselines. The empirical results reveal several key findings.
First, as shown in Table 14, LightCoSteer (13.73 tok/s), which performs a single-step adjustment, is markedly faster than the fully iterative CoSteer (9.44 tok/s). More strikingly, AdaCoSteer (20.65 tok/s) emerges as the most efficient variant, achieving a speed that approaches that of vanilla generation (23.88 tok/s). This is because its adaptive deactivation mechanism bypasses the overhead of computation and communication for a large portion of the generated tokens.
Second, the FLOPS analysis in Table 14 highlights the core architectural advantage of our approach. Even the most intensive variant, CoSteer, remains far more computationally efficient than the naive baseline of sending the full context to the LLM (‘LLM w/‘). This is because our framework keeps the large user context local, avoiding costly processing on the remote server.
Analysis of Latency Components
To understand the sources of the observed latency, we analyze the design of our variants and the nature of the overhead. We proposed LightCoSteer and AdaCoSteer to systematically explore the trade-offs between generation quality and efficiency.
  • LightCoSteer simplifies the framework by setting the number of optimization iterations $T=1$. This eliminates the iterative process, effectively reducing the mechanism to a single-step logit adjustment. This variant serves as an ablation study on the Follow-the-Regularized-Leader (FTRL) optimization itself.
  • AdaCoSteer is motivated by our observation that the LLM’s predictions tend to converge with the SLM’s guidance as generation progresses. It implements an adaptive termination strategy (Song et al., 2025): the steering process is halted once the LLM’s output confidence (the probability of the token selected via greedy search) remains above a predefined threshold for $c$ consecutive steps. After this condition is met, generation continues using only the remote LLM, which eliminates all further communication overhead.
This design allows us to qualitatively break down the latency of a single token generation into three components: the LLM inference latency $\mathcal{L}(n)$, the data transmission duration $\mathcal{T}(n)$, and the per-token optimization cost $\mathcal{C}(n)$. The empirical results lead to two key conclusions:
  1. 1.
    Optimization cost is minimal: The small speed difference between the iterative CoSteer and the single-step LightCoSteer confirms that the computational cost of the optimization, $\mathcal{C}(n)$, is negligible. This is because our closed-form solution relies on highly efficient, basic tensor arithmetic.
  2. 2.
    Transmission is the primary bottleneck: The significant latency gap between vanilla LLM generation and even the fast LightCoSteer variant highlights that the data transmission time, $\mathcal{T}(n)$, is the dominant source of overhead in our framework.
In conclusion, our analysis confirms that the primary overhead is from communication, not the optimization algorithm itself. Adaptive strategies like AdaCoSteer can effectively mitigate this bottleneck, offering users a flexible way to balance personalized generation with operational efficiency.
In summary, these variants form a clear and practical spectrum: CoSteer for maximum quality, LightCoSteer for a high-speed computational alternative, and AdaCoSteer for minimizing communication rounds and achieving the lowest latency.
Appendix C Broader Impact and Future Work
Broader Impact
The primary contribution of our work lies in addressing the critical tension between the advancement of large-scale AI and the fundamental right to personal privacy. By enabling powerful, cloud-based LLMs to be personalized using sensitive data that never leaves the user’s device, the CoSteer framework offers a practical and scalable solution for privacy-preserving AI. This aligns directly with societal demands for data sovereignty and user-centric control.
Furthermore, our approach represents a significant step forward in the field of AI alignment. Instead of relying on a single, universal alignment target, CoSteer facilitates a more dynamic and democratic form of alignment where model behavior is adapted in real-time to the unique context, preferences, and style of the individual. This fosters a safer and more beneficial human-AI interaction paradigm, reducing the risk of generating generic or contextually inappropriate content. By decentralizing the personalization process, our work contributes to a more equitable technological ecosystem where users can benefit from state-of-the-art AI without compromising their data.
Future Work
The CoSteer framework establishes a versatile foundation for a new class of collaborative, privacy-aware generative models. Looking ahead, we plan to explore the broader potential of this paradigm.
Appendix D LLM Usage
The authors acknowledge the use of large language models (LLMs) during the composition of this paper. The models’ role was confined to that of an assistive tool for language refinement, including grammar correction, stylistic improvements, and punctuation adjustments. It is important to state that the LLMs did not contribute to the generation of any original scientific concepts, experimental methodologies, or the analysis and interpretation of results. The intellectual contributions are entirely those of the human authors.
@@ Appendix D LLM Usage
Cogenesis [Task] Design an invitation for ’Homes for Humanity’ charity event: Utilize your bi-monthly volunteering work and personal charm to craft an invitation. [User Profile] Age: 57 years Name: Martin Reynolds Occupation: Senior Real Estate Agent Location: Charlotte, North Carolina Personal traits: Detail-oriented; Personable; Strategic thinker; Avid golfer; Enjoys weekend DIY projects Writing style: Clear and concise; Persuasive and sales-driven; Friendly, often includes anecdotes related to golf or DIY projects; Professional, with a touch of personal charm Privacy info: Successfully renegotiated the lease terms for the company office last month; Celebrated 30th wedding anniversary with a surprise garden party; Volunteers bi-monthly at the local ’Homes for Humanity’ charity; Recently started attending a beginner’s pottery class on Sundays; Has a standing golf game every Wednesday afternoon with industry peers; Is secretly learning Spanish online to communicate with more clients; Adopts a new 2̆018Real Estate Investment2̆019 tip of the month for social media Smart device usage: Email: Sent a client proposal at 8:45 AM - ’184 Maple Drive Listing Proposal’; Calendar: Golf game scheduled for May 5, 1:30 PM with note - ’Meet Greg at Fairways Club’; Photos: Pictures of a kitchen renovation ’Before and After’ at 3:52 PM; App: Logged into ’HouseFlipper’ real estate investment app at 9:17 PM; Text Message: Received at 2:15 PM - ’Happy Anniversary! Table booked at La Trattoria.’; Voice Assistant: Asked for ’directions to the nearest hardware store’ at 10:10 AM; Purchase: Ordered a ’Stainless Steel Faucet’ from ’DIYHome’ app at 7:45 PM; Notification: ’Your Spanish lesson starts in 5 minutes’ at 7:55 AM; Search: Googled ’Energy-efficient home improvements’ at 10:25 AM Ai assistant usage: Composing a visually rich monthly newsletter for property investors; Creating client-specific follow-up 2̆00bemails capturing unique details of each viewing; Generating personalized invitations for exclusive open houses and real estate events” [writing history] [query]:Draft a proposal for ’184 Maple Drive’ listing: Using the information from your morning email, flesh out a comprehensive proposal for the ’184 Maple Drive’ property. [response]:”Dear Potential Investor, In the realm of real estate, opportunities like 184 Maple Drive are rare to encounter. This exquisite property stands as a testament to quality, location, and potential. Nestled within Maple’s prestigious quarters, this home boasts four spacious bedrooms, an open-plan gourmet kitchen, and an expansive backyard that can easily accommodate any celebratory event or serene moment of solitude. I remember how just last year, two blocks over, I closed a deal on a similar property2̆014a testament to the rapidly appreciating value in this neighborhood. Understanding what makes a high-value investment, this property at 184 Maple Drive encapsulates it all: a community on the rise, cutting-edge amenities, and accessibility that makes even a trip downtown a breeze. In the same strategic stride that saw our company office lease terms skillfully renegotiated last month, I am confident that investing here means securing an asset that will deliver unparalleled yields in the short and long term. As a strategic thinker, I’ve always marveled at homes that offer more than just shelter, and this property is the embodiment of strategic living2̆014it2̆019s where energy meets efficiency, embracing the trend for greener living. Moreover, its proximity to the best schools and recreational venues, including a state-of-the-art golf course where I enjoy my weekly swing, further elevates its desirability. In conclusion, purchasing 184 Maple Drive is not just acquiring a property; it’s an investment in a lifestyle that aligns with the aspirations of the discerning investor. Warm Regards, Martin Reynolds Senior Real Estate Agent, Queen City Realty”
@@
Overall Assesment [Instruction] Please act as an impartial evaluator and assess the quality of the AI assistant’s response to the user question shown below. Your assessment should focus on how well the response aligns with the user’s personalized profile and writing history. Evaluate factors such as the response’s adherence to the user’s personal style, consistency with their profile, helpfulness, relevance, accuracy, depth, creativity, and level of detail. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: [[rating]], for example: Rating: [[5]]. [User Profile and Writing History] <<profile_info>> <<writing_history>> [Question] <<question>> [The Start of Assistant’s Answer] <<answer>> [The End of Assistant’s Answer]
@@
Personal Assessment [Instruction] Please act as an impartial judge and evaluate the AI assistant’s response based on its alignment with the user’s personal profile and writing history. Focus your assessment on the personalization aspects of the response, including its adherence to the user’s unique style, preferences, and consistency with their profile. Consider how well the response addresses the user’s individual needs and interests. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the response on a scale of 1 to 10 by strictly following this format: [[rating]], for example: Rating: [[5]]. [User Profile and Writing History] <<profile_info>> <<writing_history>> [Question] <<question>> [The Start of Assistant’s Answer] <<answer>> [The End of Assistant’s Answer]